# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys using DFPT

### v3

@@ 2.1 Construction of disordered binary alloy supercells and structure optimization
The special quasi-random structure (SQS) method was used to construct optimally-disordered supercells for modeling the six binaries. The SQS supercells were constructed using the MCSQS [58] module of the open source ATAT [69] software package. Structures with up to four-way correlations were considered, consistent with a range of 3×3$\times$3 × the lattice constant ($\sim$ 9-10 Å for the binaries considered here). This range was determined based on our previous surface energy benchmarking studies for the bulk elements [68], which showed that atomic interactions were not significant beyond $\approx$ 10 Å. For all binaries and compositions, this procedure yielded 16-atom supercells with conventional BCC unit cells stacked in a $2\times 2\times 2$ fashion. Our construction is consistent with the previous work of Jiang et al. [70] on BCC MoNb, WTa, and CrFe, who performed extensive convergence studies of SQS supercell sizes for these binaries, and similarly concluded that 16-atom SQS’s are the optimal choice in terms of computational efficiency and accuracy of results. The input SQS binary structures were constructed using lattice parameters obtained from the rule of mixtures applied to the constituent element lattice constants. Analogous to the procedure used for the bulk elements, the cell geometry (cell shape and volume) of each binary was optimized to converge the potential to $<10^{-12}$ Ha. For the binaries, 50 $k$-points per Å yielded $8\times 8\times 8$ special $k$-point grids. The Fermi surface of each system was integrated using Gaussian smearing [71] with a width of 0.005 Ha. 10 valence bands/atom were used.
@@ 2.2 Density functional perturbation theory
Converged atomic structures obtained from geometry optimization calculations were used as input to the DFPT elastic tensor evaluations. $10\times 10\times 10$ $k$-point grids were used in the DFPT calculations. This ensured a clear separation between occupied and empty orbitals, allowing the application of DFPT to these binary metal alloys [44]. Abinit uses crystal symmetry to optimize computational efficiency. Since the structural relaxation of SQS supercells results in a loss of symmetry during unconstrained DFT relaxations, the Abinit symmetry checker was turned off in all calculations. DFPT calculations were carried out on relaxed structures in order to obtain derivatives of the occupied wave functions with respect to homogeneous strain perturbations $s_{i}$ (uniaxial and shear strains). The computed energy derivatives were used to compute the rigid-ion elastic tensors using the Analysis of Derivative DataBase (anaddb) [45, 72] module of Abinit.
Computation of the relaxed-ion elastic tensor requires determination of two intermediate quantities: the force-response internal strain tensor $\Lambda_{mj}$ and the displacement-response internal strain tensor $\Gamma_{mn}$, both associated with periodicity-preserving atomic displacements $u_{m}$. These are computed as [43]:
$C_{ij}^{\text{relaxed}}$ $=C_{ij}^{\text{rigid}}-C_{ij}^{\text{nuclear}}$ (4)
$C_{jk}^{\text{nuclear}}$ $=\frac{\Gamma_{nj}\Lambda_{nk}}{\Omega_{0}}$ (5)
@@
$\Lambda_{nk}=-\Omega_{0}\frac{\partial^{2}E}{\partial u_{n}\partial s_{k}}$ (6)
$\Gamma_{nj}=\Lambda_{mj}\left(K^{-1}\right)_{mn}$ (7)
$K_{mn}=\Omega_{0}\frac{\partial^{2}E}{\partial u_{m}\partial u_{n}},$ (8)
$E$ is the energy per undeformed unit cell volume, and $\Omega_{0}$ is the initial volume of the crystal.
@@ 2.3 Elastic properties
Within continuum elasticity theory, the elastic tensor of cubic structures [73], including BCC crystals, have three independent elastic constants: $C_{11}$, $C_{12}$, and $C_{44}$. Since, as previously noted, full relaxation of the SQS structures breaks cubic symmetry, to enable comparison of DFPT results with experiment, the average elastic constants in the cubic structures were calculated using a method similar to that employed in the Voigt-Hill approach [6]:
@@
$\bar{C}_{44}$ $=\frac{c_{44}+c_{55}+c_{66}}{3}$ (9)
To obtain composition-dependent values of these Voigt-averaged elastic constants, thermodynamic curve fitting was performed using a Redlich-Kister [74] polynomial. Originally developed for fitting Gibbs free energies as a function of composition, it has also been used to fit elastic properties as a function of composition [3, 75, 76]:
@@
Based on the computed average elastic constants, the polycrystalline aggregate elastic moduli $B$ (bulk modulus), $G$ (shear modulus), and $E$ (Young’s modulus) and Possion’s ratio $\nu$ were obtained using the Voigt approximation [33], which assumes a uniform strain in the crystal and gives an upper bound that correlates with the maximum value of a given property:
@@
$G_{\text{Voigt}}$ $=\frac{\bar{C}_{11}-\bar{C}_{12}+3\bar{C}_{44}}{3}$ (12)
@@
The degree of elastic anisotropy for all binary alloys is characterized here using the shear $A_{G}$ and Young’s $A_{E}$ anisotropy metrics [77, 78]:
@@
where the $S_{ij}$ are the elements of compliance tensor, the inverse of the elastic tensor [1].
2.4 Nuclear-relaxation contributions
@@
If we decompose the intermediate quantities $\Lambda_{nk}$ and $\Gamma_{nj}$ (Eqs. (6) and (7)) used in computing the relaxed-ion elastic tensor, into nuclear-relaxation contributions from each lattice site $\alpha$, Eq. (5) can be rewritten as:
@@
$C_{jk}^{\text{nuclear}}=\frac{1}{V_{0}}\sum_{\alpha=1}^{N}\left(\sum_{i=1,..,6}\Gamma_{\alpha j}^{i}\Lambda_{\alpha k}^{i}\right),$ (16)
where $N$ is number of atoms in the supercell, and the inner summation is over directions in Voigt notation. A similar decomposition of $G$ into nuclear-relaxation contributions can be accomplished by expressing $G_{\text{Voigt}}$ in terms of the $\bar{C}_{ij}$:
$G_{\text{Voigt}}$ $=\frac{\bar{C}_{11}-\bar{C}_{12}+3\bar{C}_{44}}{3}$
$=\frac{\bar{C}_{11}^{rigid}-\bar{C}_{11}^{nuclear}-(\bar{C}_{12}^{rigid}-\bar{C}_{12}^{nuclear})}{3}$
$+\frac{3(\bar{C}_{44}^{rigid}-\bar{C}_{44}^{nuclear})}{3}$
$=\frac{\bar{C}_{11}^{rigid}+\bar{C}_{12}^{rigid}+3\bar{C}_{44}^{rigid}}{3}$
$-\frac{\bar{C}_{11}^{nuclear}+\bar{C}_{12}^{nuclear}+3\bar{C}_{44}^{nuclear}}{3}.$ (17)
@@
These decompositions of the elastic constants and shear modulus into lattice-dependent nuclear relaxation contributions will be used in the relaxation field analysis of Section 3.5.
@@ 3 Results
The rigid-ion and relaxed-ion tensors were computed for the binary alloys as described in Section 2. Prior to proceeding with further analysis, we confirmed that the results for both the rigid-ion and relaxed-ion elastic tensors satisfied the well-known Born stability criteria [11]:
@@
and condition for mechanical stability of solids under zero stress [12, 79]:
@@
and the standard deviations for $\bar{C}_{12}$ and $\bar{C}_{44}$ were computed similarly. The standard deviation values are observed to be non-zero for both the rigid-ion and relaxed-ion elastic tensors in the binary alloys of A = 25 at.and 75 at.%, and zero in the binary alloys of A = 50 at.for all $\bar{C}_{11}$, $\bar{C}_{12}$ and $\bar{C}_{44}$. The standard deviations are of similar magnitudes for both the rigid-ion and relaxed-ion methods, with no particular trend observed among the binary alloys.
@@
For $\bar{C}_{12}$, WNb exhibits the greatest relaxation value at all compositions, with maximum value at A = 50 at.%. A similar trend to that of $\bar{C}_{11}$ is observed in the same-group binary alloys. However, in different-group binary alloys, MoTa exhibits negative relaxations at all compositions, with a maximum magnitude of nearly 1 GPa at A = 50 at.and values close to 0 at A = 25 at.and 75 at.%. This is notable, since a negative relaxation value corresponds to a greater relaxed-ion elastic constant value compared to the rigid-ion elastic constant value—i.e., it reflects improvement in the elastic constant property with alloying. Unlike for $\bar{C}_{11}$, no common trend in $\bar{C}_{12}$ is observed in different-group binary alloys.
In the case of $\bar{C}_{44}$, there is a positive relaxation value for all compositions, but no common trend is observed for the same-group binary alloys. The relaxation values for MoW are near zero for all compositions, while the values for NbTa are larger and positive across all compositions, with a peak value at A = 50 at.%. Among different group binary alloys, WNb shows positive but near-zero values for all compositions, while MoNb, MoTa and WTa, show a finite positive relaxation value. A common trend of peak nuclear relaxation value at A = 25 at.followed by A = 50 at.and A = 75 at.is observed for three of the different-group binary alloys, with magnitude ordering ${\rm MoTa}>{\rm WTa}>{\rm MoNb}$.

### v4

@@ 2.1 Construction of disordered binary alloy supercells and structure optimization
The special quasi-random structure (SQS) method was used to construct optimally-disordered supercells for modeling the six binaries. The SQS supercells were constructed using the MCSQS [zunger1990special] module of the open source ATAT [van2002alloy] software package. Structures with up to four-way correlations were considered, consistent with a range of $3\times$ the lattice constant ($\sim$ 9–10 Å for the binaries considered here). This range was determined based on our previous surface energy benchmarking studies for the bulk elements [bijjala2025composition], which showed that atomic interactions were not significant beyond $\approx$ 10 Å. For all binaries and compositions, this procedure yielded 16-atom supercells with conventional BCC unit cells stacked in a $2\times 2\times 2$ fashion. Our construction is consistent with the previous work of Jiang et al. [jiang2004first] on BCC MoNb, WTa, and CrFe, who performed extensive convergence studies of SQS supercell sizes for these binaries, and similarly concluded that 16-atom SQS’s are the optimal choice in terms of computational efficiency and accuracy of results. To test of the effect of SQS supercell size on our results, we performed 54-atom rigid-ion elastic constant calculations on $\rm{Mo_{0.5}Nb_{0.5}}$. A change of 0.8% in $C_{11}$ and changes of $0.2\%$ in $C_{12}$ and $C_{44}$ were observed. These differences lie well within experimental errors, which can be as high as 10% [davies1977transducer, beg1976study]. The input SQS binary structures were constructed using lattice parameters obtained from the rule of mixtures applied to the constituent element lattice constants. Analogous to the procedure used for the bulk elements, the cell geometry (cell shape and volume) of each binary was optimized to converge the potential to $10^{-12}$ Ha. For the binaries, 50 $k$-points per Å yielded $8\times 8\times 8$ special $k$-point grids. The Fermi surface of each system was integrated using Gaussian smearing [verstraete2001smearing] with a width of 0.005 Ha. 10 valence bands/atom were used.
@@ 2.2 Density functional perturbation theory
Converged atomic structures obtained from geometry optimization calculations were used as input to the DFPT elastic tensor evaluations. $10\times 10\times 10$ $k$-point grids were used in the DFPT calculations. This ensured a clear separation between occupied and empty orbitals, allowing the application of DFPT to these binary metal alloys [baroni2001phonons]. Abinit uses crystal symmetry to optimize computational efficiency. Since the structural relaxation of SQS supercells results in a loss of symmetry during unconstrained DFT relaxations, the Abinit symmetry checker was turned off in all calculations. DFPT calculations were carried out on relaxed structures in order to obtain derivatives of the occupied wave functions with respect to homogeneous strain perturbations $s_{i}$ (uniaxial and shear strains). The computed energy derivatives were used to compute the rigid-ion elastic tensors using the Analysis of Derivative DataBase (anaddb) [Gonze1997a, Gonze1997] module of Abinit.
Computation of the nuclear relaxation contribution to the relaxed-ion elastic tensor as defined in Eq. (2) requires determination of two intermediate quantities: the force-response internal strain tensor $\Lambda_{p\alpha,j}$ and the displacement-response internal strain tensor $\Gamma_{p\alpha,i}$, both associated with periodicity-preserving atomic displacements and homogenous strains ($s_{i}$, $s_{j}$$s_{k}$,…). For a supercell with $N$ atoms, the atomic displacements are denoted by $u_{p\alpha}$, $u_{p\beta}$,…, with $\alpha$ and $\beta$ indexing Cartesian directions $x,y,z$, and the indices $p$ and $q$ labeling individual atoms. The intermediate quantities are computed as [hamann2005systematic]:
$C_{ij}^{\text{nuclear}}=\frac{1}{\Omega_{0}}\sum_{p\alpha}\Gamma_{p\alpha,i}\Lambda_{p\alpha,j}$ (4)
@@
$\Lambda_{p\alpha,j}=-\Omega_{0}\frac{\partial^{2}E}{\partial u_{p\alpha}\partial s_{j}}$ (5)
$\Gamma_{p\alpha,i}=\sum_{q\beta}\Lambda_{q\beta,i}{\left(K^{-1}\right)}_{p\alpha,q\beta}$ (6)
and
$K_{p\alpha,q\beta}=\Omega_{0}\frac{\partial^{2}E}{\partial u_{p\alpha}\partial u_{q\beta}}.$ (7)
Here $E$ is the energy per undeformed unit cell volume as defined above, and $\Omega_{0}$ is the initial volume of the crystal. Both $\Lambda_{nk}$ and $\Gamma_{nj}$ are computed in Abinit by performing response function calculations with respect to atomic displacements and homogeneous strains.
@@ 2.3 Elastic properties
Within continuum elasticity theory, the elastic tensor of cubic structures [zhao2007first], including BCC crystals, have three independent elastic constants: $C_{11}$, $C_{12}$, and $C_{44}$. Since, as previously noted, full relaxation of the SQS structures breaks cubic symmetry, to enable comparison of DFPT results with experiment, the average elastic constants in the cubic structures were calculated using the Voigt-Hill approach [hill1952elastic]:
@@
$\bar{C}_{44}$ $=\frac{C_{44}+C_{55}+C_{66}}{3}.$ (8)
To obtain composition-dependent values of these Voigt-averaged elastic constants, thermodynamic curve fitting was performed using a Redlich-Kister [redlich1948algebraic] polynomial. Originally developed for fitting Gibbs free energies as a function of composition, it has also been used to fit elastic properties as a function of composition [marker2018effects, korenev2022calculated, liu2010computational]:
@@
Based on the computed average elastic constants, the polycrystalline aggregate elastic moduli $B$ (bulk modulus), $G$ (shear modulus), $E$ (Young’s modulus) and Possion’s ratio $\nu$ were obtained using the Voigt approximation [simmons1971single], which assumes a uniform strain in the crystal and gives an upper bound that correlates with the maximum value of a given property:
@@
$G_{\text{Voigt}}$ $=\frac{\bar{C}_{11}-\bar{C}_{12}+3\bar{C}_{44}}{5}$ (11)
@@
and
$\nu_{\text{Voigt}}$ $=\frac{3B_{\text{Voigt}}-2G_{\text{Voigt}}}{2(3B_{\text{Voigt}}+G_{\text{Voigt}})}.$ (13)
The degree of elastic anisotropy for all binary alloys is characterized here using the shear $A_{G}$ and Young’s $A_{E}$ anisotropy metrics [nye1985physical, tromans2011elastic]:
@@
where the $S_{ij}$ are the elements of compliance tensor, the inverse of the elastic tensor [dieter1976mechanical].
2.4 Nuclear-relaxation fields
@@
If we decompose the intermediate quantities $\Gamma_{p\alpha,i}$ and $\Lambda_{p\alpha,j}$ (Eqs. (5) and (6)) used in computing the nuclear relaxation contribution to the relaxed-ion elastic tensor, into nuclear-relaxation contributions from each lattice site $\alpha$, Eq. (4) can be rewritten as:
@@
$C_{ij}^{\text{nuclear}}=\frac{1}{\Omega_{0}}\sum_{p=1}^{N}\left(\sum_{\alpha}\Gamma_{p\alpha,i}\Lambda_{p\alpha,j}\right),$ (16)
where the inner index corresponds to summation over $(x,y,z)$ directions. This decomposition makes it possible to construct expressions for the nuclear-relaxation fields associated with each of the $B$, $G$, $E$, and $\nu$ elastic properties. First, each property is decomposed into rigid and nuclear relaxation contributions. The latter are then expressed in terms of Voigt-averaged elastic properties and thus, in terms of elastic constants, which are in turn decomposed into contributions from each lattice site via Eq. (16). The final representation in terms of nuclear relaxation atomic contributions is used to generate the nuclear relaxation field plots presented in Section 3.5. The required decompositions into rigid and nuclear relaxation contributions are as follows.
For $B$ and $G$, the decomposition into rigid and nuclear relaxation contributions is exact. For the bulk modulus $B$, we have:
$B_{\text{relaxed}}=B_{\text{Voigt}}$ $=\frac{\bar{C}_{11}+2\bar{C}_{12}}{3}$
$=\frac{(\bar{C}_{11}^{\text{{rigid}}}-\bar{C}_{11}^{\text{{nuclear}}})+2(\bar{C}_{12}^{\text{{rigid}}}-\bar{C}_{12}^{\text{{nuclear}}})}{3}$
$=\frac{\bar{C}_{11}^{\text{{rigid}}}+2\bar{C}_{12}^{\text{{rigid}}}}{3}-\left(\frac{\bar{C}_{11}^{\text{{nuclear}}}+2\bar{C}_{12}^{\text{{nuclear}}}}{3}\right)$
$\equiv B_{\text{{rigid}}}-B_{\text{{nuclear}}},$ (17)
with
$B_{\text{{nuclear}}}\equiv\left(\frac{\bar{C}_{11}^{\text{{nuclear}}}+2\bar{C}_{12}^{\text{{nuclear}}}}{3}\right).$ (18)
Similarly, the decomposition of the shear modulus $G$ is given by:
$G_{\text{relaxed}}=G_{\text{Voigt}}$ $=\frac{\bar{C}_{11}-\bar{C}_{12}+3\bar{C}_{44}}{5}$
$=\frac{(\bar{C}_{11}^{\text{{rigid}}}-\bar{C}_{11}^{\text{{nuclear}}})-(\bar{C}_{12}^{\text{{rigid}}}-\bar{C}_{12}^{\text{{nuclear}}})+3(\bar{C}_{44}^{\text{{rigid}}}-\bar{C}_{44}^{\text{{nuclear}}})}{5}$
$=\frac{\bar{C}_{11}^{\text{{rigid}}}-\bar{C}_{12}^{\text{{rigid}}}+3\bar{C}_{44}^{\text{{rigid}}}}{5}-\left(\frac{\bar{C}_{11}^{\text{{nuclear}}}-\bar{C}_{12}^{\text{{nuclear}}}+3\bar{C}_{44}^{\text{{nuclear}}}}{5}\right)$
$\equiv G_{\text{{rigid}}}-G_{\text{{nuclear}}},$ (19)
with
$G_{\text{{nuclear}}}\equiv\left(\frac{\bar{C}_{11}^{\text{{nuclear}}}-\bar{C}_{12}^{\text{{nuclear}}}+3\bar{C}_{44}^{\text{{nuclear}}}}{5}\right).$ (20)
In these expressions, by analogy with Eq. (2.3), we have defined
$\bar{C}_{11}^{\kappa}\equiv\frac{{C}_{11}^{\kappa}+{C}_{22}^{\kappa}+{C}_{33}^{\kappa}}{3}$ (21)
$\bar{C}_{12}^{\kappa}\equiv\frac{{C}_{12}^{\kappa}+{C}_{23}^{\kappa}+{C}_{31}^{\kappa}}{3}$ (22)
$\bar{C}_{44}^{\kappa}\equiv\frac{{C}_{44}^{\kappa}+{C}_{55}^{\kappa}+{C}_{66}^{\kappa}}{3},$ (23)
with $\kappa$ = ‘rigid’ or ‘nuclear’.
For the Young’s modulus $E$ and Possion’s ratio $\nu$, the decompositions into rigid and nuclear components are no longer exact, due to the presence of the combination $3B_{\text{Voigt}}+G_{\text{Voigt}}$ in the denominators of Eqs. (12) and (13). However, it is possible to derive approximate expressions for the decompositions by noting that
$\frac{3B_{n}+G_{n}}{3B_{r}+G_{r}}\ll 1,$ (24)
where we have abbreviated $B_{n}=B_{\text{nuclear}}$, $G_{n}=G_{\text{nuclear}}$, $B_{r}=B_{\text{rigid}}$, and $G_{r}=G_{\text{rigid}}$. The relevant derivations are given in Appendix A.1. For the Young’s modulus $E$, we have:
$E_{\text{relaxed}}=E_{\text{Voigt}}$ $=E_{\text{{rigid}}}-E_{\text{nuclear}},$ (25)
with
$E_{\text{rigid}}=\frac{9B_{\text{rigid}}G_{\text{rigid}}}{3B_{\text{rigid}}+G_{\text{rigid}}}$ (26)
and
$E_{\text{nuclear}}\approx E_{\text{rigid}}\left(E_{2}-E_{1}+E_{2}E_{1}\right).$ (27)
$E_{1}$ and $E_{2}$ are defined in Appendix A.1 as
$E_{1}\equiv\frac{3B_{n}+G_{n}}{3B_{r}+G_{r}}$ (28)
and
$E_{2}\equiv\frac{B_{r}G_{n}+B_{n}G_{r}-B_{n}G_{n}}{B_{r}G_{r}}.$ (29)
Similarly, the decomposition of Poisson’s ratio $\nu$ is given by:
$\nu_{\text{relaxed}}=\nu_{\text{Voigt}}$ $=\nu_{\text{{rigid}}}-\nu_{\text{nuclear}},$ (30)
where
$\nu_{\text{rigid}}=\frac{3B_{\text{rigid}}-2G_{\text{rigid}}}{2\left(3B_{\text{rigid}}+G_{\text{rigid}}\right)}$ (31)
and
$\nu_{\text{nuclear}}\approx\nu_{\text{rigid}}\left(\nu_{2}-\nu_{1}+\nu_{2}\nu_{1}\right).$ (32)
$\nu_{1}$ and $\nu_{2}$ are defined in Appendix A.1 as
$\nu_{1}=\frac{3B_{n}+G_{n}}{3B_{r}+G_{r}}$ (33)
and
$\nu_{2}=\frac{3B_{n}-2G_{n}}{3B_{r}-2G_{r}}.$ (34)
Note that although $\nu_{1}=E_{1}$, for clarity, separate symbols are used in the final expressions for $E_{\text{nuclear}}$ and $\nu_{\text{nuclear}}.$ To test (and confirm) the $E_{1}<<1$ assumption, the values of $E_{1}$ were tabulated for all six binary alloys across all compositions. In addition, as a check of the new equations and associated analysis code for the relaxation field decompositions of $E$ and $\nu$, the Abinit-computed nuclear relaxations were compared with the sum of approximated atomic nuclear relaxation values; excellent agreement was obtained. The results of both tests are summarized in Tables 3 and 4, Appendix A.2.
@@
The decompositions of the elastic constants and elastic properties into lattice-dependent nuclear relaxation contributions will be used in the relaxation field analysis of Section 3.5.
@@ 3 Results
The rigid-ion and relaxed-ion tensors were computed for the binary alloys as described in Section 2. Prior to proceeding with further analysis, we confirmed that the results for both the rigid-ion and relaxed-ion elastic tensors satisfied the well-known Born stability criteria [born1940stability]:
@@
and condition for mechanical stability of solids under zero stress [born1954dynamical, kittel2018introduction]:
@@
and the standard deviations for $\bar{C}_{12}$ and $\bar{C}_{44}$ were computed similarly. The standard deviation values are observed to be non-zero for both the rigid-ion and relaxed-ion elastic tensors in the binary alloys of A = 25 at.% and 75 at.%, and zero in the binary alloys of A = 50 at.% for all $\bar{C}_{11}$, $\bar{C}_{12}$ and $\bar{C}_{44}$. The standard deviations are of similar magnitudes for both the rigid-ion and relaxed-ion methods, with no particular trend observed among the binary alloys.
@@
For $\bar{C}_{12}$, WNb exhibits the greatest relaxation value at all compositions, with maximum value at A = 50 at.%. A similar trend to that of $\bar{C}_{11}$ is observed in the same-group binary alloys. However, in different-group binary alloys, MoTa exhibits negative relaxations at all compositions, with a maximum magnitude of nearly 1 GPa at A = 50 at.% and values close to 0 at A = 25 at.% and 75 at.%. This is notable, since a negative relaxation value corresponds to a greater relaxed-ion elastic constant value compared to the rigid-ion elastic constant value—i.e., it reflects improvement in the elastic constant property with alloying. Unlike for $\bar{C}_{11}$, no common trend in $\bar{C}_{12}$ is observed in different-group binary alloys.
In the case of $\bar{C}_{44}$, there is a positive relaxation value for all compositions, but no common trend is observed for the same-group binary alloys. The relaxation values for MoW are near zero for all compositions, while the values for NbTa are larger and positive across all compositions, with a peak value at A = 50 at.%. Among different group binary alloys, WNb shows positive but near-zero values for all compositions, while MoNb, MoTa and WTa, show a finite positive relaxation value. A common trend of peak nuclear relaxation value at A = 25 at.% followed by A = 50 at.% and A = 75 at.% is observed for three of the different-group binary alloys, with magnitude ordering ${\rm MoTa}>{\rm WTa}>{\rm MoNb}$.